# Pretraining and Updating Language- and Domain-specific Large Language Model: A Case Study in Japanese Business Domain

### v2

@@ Abstract
Several previous studies have considered language- and domain-specific large language models (LLMs) as separate topics. This study explores the combination of a non-English language and a high-demand industry domain, focusing on a Japanese business-specific LLM. This type of a model requires expertise in the business domain, strong language skills, and regular updates of its knowledge. We trained a 13-billion-parameter LLM from scratch using a new dataset of business texts and patents, and continually pretrained it with the latest business documents. Further we propose a new benchmark for Japanese business domain question answering (QA) and evaluate our models on it. The results show that our pretrained model improves QA accuracy without losing general knowledge, and that continual pretraining enhances adaptation to new information. Our pretrained model and business domain benchmark are publicly available.  [^1] https://huggingface.co/stockmark
@@ 1 Introduction
This study explores a language- and domain-specific large language model (LLM), a Japanese LLM for the business domain. As examples of business-related questions shown in Table 1, answering these questions requires knowledge on current affairs, corporate activities, and/or social issues. Although LLMs are successfully applied to various language tasks, models are pretrained on datasets consisting primarily of general English data. Consequently, their performance in other languages, such as Japanese, remains underexplored, particularly in domain-specific tasks.
Our study aligns with two existing directions: language-specific and domain-specific LLMs. Language-specific pretrained models have been developed for various languages (Zhang et al., 2021; dbmdz, 2023; tokyotech llm, 2023) nowadays, whereas domain-specific models excel in several domains such as finance (Scao et al., 2023). The combination of these two aspects with language models have been known promising since the era of BERTology (Beltagy et al., 2019; Ishigaki et al., 2023), however, many non-English languages, including Japanese, have no such LLM combinations for the recent decoder-only architecture such as GPTs (Brown et al., 2020). In this situation, first and foremost, a fundamental decoder-only model is required to facilitate future research to discover better pretraining methods. Therefore, as a case study, we developed the first Japanese LLM with 13 billion parameters for the business domain.
To leverage our Japanese business specific LLM in practice, it should be updated, for example, to answer the latest business-related questions. Thus, we adopt a pipeline approach for training our LLM: pretraining and continual updating. For the pretraining step, we train the model from scratch. Although there are two potential approaches – pretraining from scratch and continual pretraining from an existing LLM – we opt for the former because ample data are available in the target domain and language.
For the training dataset, we compile a diverse set that combines domain-specific and general-domain data. Specifically, we use over 40 billion tokens from patents and business documents to capture technical terminology, along with over 150 billion tokens from general-domain sources, such as Wikipedia, CC100, mC4, and Common Crawl.
To update our model with the latest business knowledge, we continually pretrain the model on the latest business documents collected in the past two months, ensuring that its knowledge remains current.
In addition to the commonly used general-domain benchmark such as ‘lm-evaluation-harness‘ (Gao et al., 2023), a business-specific benchmark is required to evaluate our business-specific model. For this purpose, we design a new dataset of 50 business questions for three different QA tasks: question-only, question with automatically retrieved context, and question with manually retrieved context. Manual human evaluations of the domain-specific benchmark demonstrate that our pretrained models outperform existing Japanese LLMs in terms of accuracy, particularly in the question-only setting. Moreover, we find that the updated models with the latest knowledge are more accurate than the pretrained models in answering questions about the most recent events in the last two months.
Our contributions are threefold: 1) we pretrain the first and largest Japanese domain-specific LLM on a new dataset; 2) we create a new benchmark with 50 questions for three different tasks; and 3) we demonstrate the effectiveness of our pretrained model in domain-specific tasks. This study establishes a fundamental environment for comparing future language- and domain-speciifc LLMs. Analysis presented in this paper provides insights for researchers who train domain- and language-specific LLMs in other domains and languages. Our model and benchmark questions are publicly available.  [^2] https://huggingface.co/stockmark.
@@
Which country joined NATO in April 2023 in response to
@@
Which Japanese startup has been developing perovskite
@@ 2 Methods
Our approach consists of three steps: collecting datasets, filtering them, and pretraining a LLM from scratch.
@@ 2.1 Dataset
To construct our pretraining dataset, we gathered 19.8% domain-specific and 80.2% general domain texts in Japanese. For the domain-specific data, we use business-related web pages and patent documents. The former, called the “Curated Business Corpus,” comprises publicly available web pages published by September 2023. We identified pages related to business using predefined URLs and a cue word list, and extracted the pages of which match the pattern of the URL, or which contain at least one of these keywords. The URLs and cue words are chosen to cover various aspects of the business domain, including chemistry, materials, biology, engineering, economics, current affairs, and social trends. In addition, we incorporate patent documents from the Japan Patent Office to enhance the model’s understanding of technical terminology. We collected as many articles as possible from these sources.
To ensure diversity of the training data, we include broader general-domain data from sources such as Wikipedia, CC100, mC4, and Common Crawl. These datasets contribute to general knowledge, which is essential for diverse natural language processing tasks.
@@ 2.2 Filtering
Filtering is critical for enhancing the dataset quality. We applied a three-step pipeline: language identification, removal of noisy characters, and deduplication.
Language Identification: Language identification is an important step in language-specific settings. To identify non-Japanese documents, we used a pipeline of two methods: the library-based and language characteristics-based. We initially used the “xlm-roberta-base-language-detection” library  [^3] https://huggingface.co/papluca/xlm-roberta-base-language-detection. Consequently, for documents with uncertain language identification results from the xlm-roberta-based tool, we applied the “franc” library  [^4] https://github.com/wooorm/franc. Any texts not recognized as Japanese were subsequently removed from the dataset.
Noise Character Removal: We believe that low-quality Japanese texts often take the non-sentential forms. Noisy texts are often composed only of dates, menu bars, URLs, or do not contain an end-of-sentence punctuation mark, “。” (period in Japanese.). Thus, we removed those texts from the corpus if they are regarded as non-sentential. This ensures the quality of our dataset and focuses on the model’s attention to relevant linguistic features. Also, because English sentences end with a period (“.”), using puctuation marks as a clue to identify non-sentential texts can aid in refining English sentences. Our characteristics-based method is simple but it has wide applicability in distinguishing languages that have specific characteristics from English; for example, Thai and Chinese do not have punctuation.
Deduplication: Finally, we deduplicated the collected documents and sentences to remove identical entries. For the document-level, we used “hash“, which is a built-in function of python, to convert documents into hashed values and removed duplicates memory-efficiently. For the sentence-level, we count the times of the sentence appearance and removed the sentences that appear more than 15 times. In both, document- an sentence-wise deduplication, we used exact matching of entries.
Following these preprocesses, our dataset comprised a total of 220 billion tokens, as shown in Table 2.
@@ 2.3 Pretraining
We use the same architecture as Llama2 (Touvron et al., 2023) with 13 billion parameters. We aligned our hyperparameters with those reported in the paper of Llama2.
For the training data in an epoch, we adopted the weighting strategy of the data used by Llama2, and doubled the amount of Wikipedia for two main reasons. First, Wikipedia data tends to be relatively clean. Second, Wikipedia contains a wealth of contents related to the business domain. In addition, we also double the data size of our Curated Business Corpus to enhance the infusion of domain-specific knowledge. By increasing the proportion of clean domain-specific data, we aim to mitigate the impact of noisy data sources such as mC4 and Common Crawl.
For the infrastructure, we used AWS’s Trainium, a hardware accelerator designed specifically for high-performance machine learning computations. We deployed our scripts on 16 trn1.32xlarge instances, each equipped with Trainium, to facilitate a robust distributed learning environment. The distributed learning process on Trainium was managed using the neuronx-nemo-megatron library  [^5] https://github.com/aws-neuron/neuronx-nemo-megatron, which is available on AWS’s custom accelerators. The pretraining phase spanned 30 days to complete training one epoch of the data.
Dataset Number of tokens [billion]
@@
Dataset Number of examples language
@@
2.4 Updating Pretrained Model with Latest Business Documents
In real-world applications, such as domain-specific QA, LLMs must be updated using the latest knowledge. However, the pretrained model discussed in the previous sections is trained on texts published only up to September 2023. It may be difficult to answer questions regarding events after October 2023.
To address this, we continue to train our pretrained model using the latest business documents published in October and November 2023 in our experiments. However, we are aware of the phenomenon of “catastrophic forgetting,” which occurs when acquiring new knowledge and potentially displaces previously learned information during pretraining.
To prevent this, we use a strategy that blends the latest business documents with randomly selected non-latest documents from the Curated Business Corpus. This strategy is inspired by previous research (Scialom et al., 2022). We introduce a hyperparameter $r$, which represents the proportion of the instances sampled from the set of non-latest documents; that is, we used the Curated Business Corpus as the source. For example, if $r$ is set to $0.3$, it means that 30% of the data of the continual update comes from the Curated Business Corpus. We investigate how different values of $r$ influence the performance of the domain-specific QA, and analyze the occurrence of catastrophic forgetting.
@@ 3 Experiments
To assess our LLMs, we create a benchmark for QA within the business domain. We then compared our models against different Japanese LLMs using this benchmark, along with the common Japanese benchmarks.
@@ 3.1 Benchmarks
Business Question Benchmark: This dataset comprises of 50 questions written in natural sentences, each paired with related articles. The questions cover various topics, including recent events, company activities, social issues, and business trends. Each question is linked to web pages retrieved through different methods: automatic and manual. The benchmark offers three QA settings: 1) answering without any context (NoContext-QA), 2) using an automatically retrieved webpage (AutoRAG-QA), and 3) using a manually retrieved webpage (ManualRAG-QA).
For the first task (NoContext-QA), web pages are not provided as a context. This setting allows us to assess the LLMs’ ability to generate answers based on the knowledge they contain internally, without knowledge existing outside the LLM.
@@
For the latter two tasks (AutoRAG-QA and ManualRAG-QA), we extract only the text from each web page and truncate it to 1000 characters.
For AutoRAG-QA, we use a search engine to identify the most relevant web pages for each question. Subsequently, we selected the highest-ranked page of which the body text was available to be extracted as the context. However, because these web pages are automatically retrieved, they may not always provide answers to questions. Thus, this setting facilitates the assessment of how well LLMs can generate an answer, considering both the knowledge they already possess and any potentially irrelevant information on the web page.
For ManualRAG-QA we manually select a web page containing the answers to the question. Note that these web pages must contain answers to the questions; thus, this task is rather a comprehension task that requires better language reading ability than the ability to produce an answer from the knowledge in LLMs. We used the prompts shown in Appendix 1.
@@
During the evaluation using this benchmark, the responses generated by the LLMs were manually assessed by an NLP researcher. The evaluation involves a binary judgement of the responses as either correct or incorrect, based on two main criteria. 1) Content faithfulness: the response must contain an answer to the question without any factual errors. 2) Response appropriateness: if the question included specific instructions (e.g. “provide only one example”), the response is considered correct only if it follows the instructions. The evaluator judges the output as correct if it satisfies both criteria. We used manual evaluations over automatic evaluation metrics because the latter (e.g., BLEU (Papineni et al., 2002), METEOR (Banerjee & Lavie, 2005), BERTScore (Zhang et al., 2020)) do not adequately assess factual correctness, which is crucial in the business domain. Moreover, if a response contains repetition, for example, repeating the same sentence twice, redundant parts were disregarded, and the evaluation focuses solely on the unique content.
@@
lm-evaluation-harness: lm-evaluation-harness  [^6] https://github.com/Stability-AI/lm-evaluation-harness is a common existing framework for assessing Japanese LLMs. It comprises eight Japanese NLP tasks in the general-domain. Each task is designed to probe different facets of language comprehension and generation. The tasks in this framework include reading comprehension (JSQuAD (Kurihara et al., 2022)), QA (JCommonSenseQA (Kurihara et al., 2022) and JAQKET  [^7] https://www.nlp.ecei.tohoku.ac.jp/projects/jaqket/), natural language inference (JNLI (Kurihara et al., 2022)), summarization (XLSum-ja  [^8] https://huggingface.co/datasets/mkshing/xlsum_ja), co-reference resolution (WinoGrande-ja (Takahashi et al., 2023)), and math problems (MGSM (Zhang et al., 2023)).
@@
Owing to the unavailability of the MARC-ja dataset (Kurihara et al., 2022), we conducted lm-evaluation-harness on the remaining seven tasks. It is well acknowledged that the efficacy of LLMs is significantly influenced by the prompts or choice of the templates used in the evaluation. To this end, because multiple templates are available for a task, we conducted evaluations across all available templates, and the highest score is reported in this paper. The configurations of templates used in this study are listed in Table 9 for reproducibility.
3.2 Compared Models and Instruction-tuning

### v3

# Pretraining and Updates of Domain-Specific LLM: A Case Study in the Japanese Business Domain

@@ Abstract
The development of Large Language Models (LLMs) in various languages has been advancing, but the combination of non-English languages with domain-specific contexts remains underexplored. This paper presents our findings from training and evaluating a Japanese business domain-specific LLM designed to better understand business-related documents, such as the news on current affairs, technical reports, and patents. Additionally, LLMs in this domain require regular updates to incorporate the most recent knowledge. Therefore, we also report our findings from the first experiments and evaluations involving updates to this LLM using the latest article data, which is an important problem setting that has not been addressed in previous research. From our experiments on a newly created benchmark dataset for question answering in the target domain, we found that (1) our pretrained model improves QA accuracy without losing general knowledge, and (2) a proper mixture of the latest and older texts in the training data for the update is necessary. Our pretrained model and business domain benchmark are publicly available  [^1] https://huggingface.co/stockmark to support further studies.
Pretraining and Updates of Domain-Specific LLM:
A Case Study in the Japanese Business Domain
Kosuke Takahashi, Takahiro Omi, Kosuke Arima Stockmark {kosuke.takahashi, takahiro.omi, kosuke.arima}@stockmark.co.jp
Tatsuya Ishigaki National Institute of Advanced Industrial Science and Technology ishigaki.tatsuya@aist.go.jp
@@ 1 Introduction
The development of Large Language Models (LLMs) has seen significant progress across various languages. However, the combination of language-specific and domain-specific contexts remains underexplored. This study focuses on a Japanese LLM tailored for the business domain, addressing the need for models that can understand and process business-related documents such as articles on current affairs, corporate activities, and social issues.
@@
Which country joined NATO in
April 2023 in response to
@@
Which Japanese startup has
been developing perovskite
@@
As demonstrated by the examples of business-related questions in Table 1, accurately answering these questions requires specialized knowledge about current events, corporate activities, and social issues. Despite the success of LLMs in various language tasks, most models are pretrained on datasets consisting primarily of general English data. Consequently, their performance in other languages, including Japanese, and in domain-specific tasks may remain limited.
Our research aligns with two key directions in LLM development: language-specific models and domain-specific models. Language-specific pretrained models have been developed for several languages (Zhang et al., 2021; dbmdz, 2023; tokyotech llm, 2023), while domain-specific models have excelled in areas such as finance (Scao et al., 2023). Combining these aspects has been shown promising since the traditional advent of domain-specific BERTs (Beltagy et al., 2019; Ishigaki et al., 2023). However, many non-English languages, including Japanese, lack such combinations, especially for recent decoder-only architectures like GPTs (Brown et al., 2020). Addressing this gap requires developing fundamental domain-specific GPT. As a case study, we present the first Japanese LLM with 13 billion parameters specifically for the business domain.
LLMs in the business domain needs to be updated regularly to address the latest business-related queries. For example, LLMs pretrained before 2023 do not know results of the Olympic games in 2024. While the importance of continual updates, there has been limited research on how we can better update LLMs with the latest knowledge without losing the general knowledge. This paper focuses on a research question about the data used for updates: how can we mix recent texts with general texts to ensure that the LLM incorporates up-to-date information while still retaining general knowledge? To address this gap, we continually pretrain our LLM with recent business documents with various ways of the mixture, ensuring that its knowledge remains current. This approach aims to improve the model’s accuracy in reflecting the latest business trends and information.
Evaluating a domain-specific model presents unique challenges. In addition to using general-domain benchmarks such as lm-evaluation-harness (Gao et al., 2023), we introduce a new business-specific benchmark. This benchmark consists of business questions across three tasks whose inputs are 1) question-only, 2) question with automatically retrieved context, and 3) question with manually retrieved context. Manual evaluations on this benchmark reveal that our pretrained model outperforms existing general-domain Japanese LLMs in accuracy, particularly in the question-only setting. Furthermore, models updated with the latest knowledge show improved accuracy in answering questions about recent events.
Our contributions are threefold: (1) we pretrain the first and largest Japanese business domain-specific LLM; (2) we present the first experiments on LLMs updated with the latest business documents and demonstrate that a proper mixture of recent and older texts in the training data is necessary; (3) we create a new benchmark with questions designed for three distinct tasks; and (4) we demonstrate the effectiveness of our pretrained model in domain-specific tasks. This study establishes a foundational environment for comparing future language- and domain-specific LLMs and provides valuable insights for researchers working with LLMs in other domains and languages. Our model and benchmark questions are publicly available  [^2] https://huggingface.co/stockmark/stockmark-13b
https://huggingface.co/datasets/stockmark/business-questions
@@ 2 Methods
Our approach consists of three steps: collecting datasets, filtering them, and pretraining an LLM from scratch. The trained LLM is later used for updating.
@@ 2.1 Dataset
To construct our pretraining dataset, we collected 19.8% domain-specific texts and 80.2% general-domain texts in Japanese.
For the domain-specific data, we created our original “Curated Business Corpus” and also used patent documents from the Japan Patent Office. These two corpora are designed to provide the business knowledge and technical terminology necessary for the target domain. The Curated Business Corpus was built by curating publicly available web pages published up to September 2023. We identified relevant pages using predefined URLs and a list of cue words, extracting pages that matched the URL patterns or contained at least one of the keywords. The URLs and cue words were selected to cover various aspects of the business domain, including chemistry, materials, biology, engineering, economics, current affairs, and social trends.
To ensure a diverse training dataset, we also included general-domain data from sources such as Wikipedia, CC100, mC4, and Common Crawl. These datasets provide essential general knowledge, which is crucial for handling a wide range of natural language processing tasks. Most other LLMs utilize similar general-domain datasets, though often with different filtering strategies.
@@ 2.2 Filtering
Filtering is essential for enhancing dataset quality. We implemented a three-step pipeline: language identification, removal of noisy characters, and deduplication.
Language Identification: Language identification is crucial for language-specific datasets. We used a two-method pipeline to identify non-Japanese documents: a library-based method and a language characteristics-based method. Initially, we employed the ”xlm-roberta-base-language-detection” library  [^3] https://huggingface.co/papluca/xlm-roberta-base-language-detection. For documents with uncertain results from this tool, we applied the franc library  [^4] https://github.com/wooorm/franc. Any texts not recognized as Japanese were removed from the dataset.
Noise Character Removal: Low-quality Japanese texts often lack proper sentence structure. Noisy texts may consist only of dates, HTML tags of menu bars, URLs, or lack end-of-sentence punctuation marks, such as “{CJK}UTF8ipxg。” (Japanese period). We removed such texts if they were deemed non-sentential to ensure the dataset’s quality and to focus the model on relevant linguistic features. Additionally, because English sentences end with a period (“.”), using punctuation as a clue helps in refining English sentences. Our characteristics-based method, while simple, is effective in distinguishing languages with specific features, such as Thai which does not use punctuation.
Deduplication: Finally, we deduplicated the collected documents and sentences to eliminate identical entries. At the document level, we used Python’s built-in hash function to convert documents into hashed values, which allowed us to remove duplicates efficiently. At the sentence level, we counted the frequency of sentences and removed those appearing more than 15 times. Both document-level and sentence-level deduplication were performed using exact matching. This step prevents the model from fixating on repeated data and ensures dataset diversity.
Following these preprocessing steps, our dataset comprised a total of 220 billion tokens, as detailed in Table 2.
@@ 2.3 Pretraining
We used the Llama2 architecture (Touvron et al., 2023) with 13 billion parameters for our model. Our hyperparameters were aligned with those reported in the Llama2 paper. Instead of further training Meta’s Llama2 weights, our model was pretrained from scratch.
To balance the training data for each epoch, we adopted the weighting strategy used by Llama2, but doubled the amount of Wikipedia data for two main reasons. First, Wikipedia data is relatively clean. Second, it contains a wealth of content relevant to the business domain. Additionally, we doubled the size of our Curated Business Corpus to enhance the integration of domain-specific knowledge. By increasing the proportion of clean, domain-specific data, we aimed to mitigate the impact of noisy data sources like mC4 and Common Crawl.
For infrastructure, we utilized AWS’s Trainium, a hardware accelerator specifically designed for high-performance machine learning computations. We deployed our training scripts on 16 trn1.32xlarge instances, each equipped with Trainium, to create a robust distributed learning environment. The distributed learning process was managed using the neuronx-nemo-megatron library  [^5] https://github.com/aws-neuron/neuronx-nemo-megatron, which is available on AWS’s custom accelerators. The pretraining phase spanned 30 days to complete one epoch of training data.
Dataset Num. of tokens [billion]
@@
Dataset Num. of examples Language
@@
2.4 Updating the Pretrained Model with Latest Business Documents
In real-world applications, such as domain-specific question answering (QA), LLMs must be updated to incorporate the most recent knowledge. The pretrained model discussed in the previous sections was trained on texts published up until September 2023. Consequently, it may struggle to answer questions about events occurring after October 2023.
To address this limitation, we continued training our pretrained model with the latest business documents published in October and November 2023. However, we are mindful of the issue of “catastrophic forgetting” (French, 1999), where acquiring new knowledge can unintentionally displace previously learned information.
To mitigate catastrophic forgetting, we employ a strategy that blends the latest business documents with randomly selected, non-latest documents from the Curated Business Corpus. This approach is inspired by previous research (Scialom et al., 2022). We introduce a hyperparameter $r$, which represents the proportion of instances sampled from the non-latest document set. For instance, if $r$ is set to 0.3, then 30% of the continual update data comes from the Curated Business Corpus.
Our experiments in the next sections show how different values of $r$ affect the performance of domain-specific QA and analyze the extent of catastrophic forgetting.
@@ 3 Experiments
To evaluate our LLM, we have created a benchmark, Business Question Benchmark, for business domain question answering. We then compared our model against various Japanese LLMs using this benchmark, as well as common Japanese benchmarks, lm-evaluation-harness.
@@ 3.1 Benchmarks
Business Question Benchmark: This benchmark consists of 50 questions written in natural language, each paired with relevant articles. The questions cover a range of topics, including recent events, company activities, social issues, and business trends. Each question is associated with web pages retrieved through automatic and manual methods.
The benchmark offers three QA settings; 1) NoContext-QA: In this setting, no web pages are provided as context. This allows us to assess the LLMs’ ability to generate answers based solely on their internal knowledge, without external information, 2) AutoRAG-QA: for this task, we use a search engine to find the most relevant web pages for each question. The highest-ranked page with available body text is selected as the context. This setting helps evaluate how well LLMs can generate answers considering both their existing knowledge and potentially irrelevant information from the web page, 3) ManualRAG-QA: Here, we manually select a web page that contains answers to the question. The RAG tasks are more about comprehension, requiring the model to understand and extract information from a specific page rather than relying solely on its internal knowledge.
@@
For AutoRAG-QA and ManualRAG-QA, the text from each web page is truncated to 1000 characters. The prompts used for these tasks are detailed in Appendix 2.
@@
Responses generated by the LLMs during the evaluation were manually assessed by an NLP researcher. The evaluation involved a binary judgment of responses as either correct or incorrect based on two main criteria: 1) Content Faithfulness: The response must accurately answer the question without any factual errors, 2) Response Appropriateness: If the question included specific instructions (e.g., “provide only one example”), the response is considered correct only if it follows these instructions.
The evaluator considered a response correct if it met both criteria. Manual evaluation was preferred over automatic metrics (e.g., BLEU (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), BERTScore (Zhang et al., 2020)) because these metrics do not adequately assess factual correctness, which is crucial in the business domain. Additionally, redundant parts of responses, such as repeated sentences, were disregarded, focusing only on unique content.
@@
lm-evaluation-harness: The lm-evaluation-harness framework  [^6] https://github.com/Stability-AI/lm-evaluation-harness is a well-established tool for evaluating Japanese LLMs. It includes eight Japanese NLP tasks spanning various domains of language comprehension and generation. These tasks include reading comprehension (JSQuAD (Kurihara et al., 2022)), QA (JCommonsenseQA (Kurihara et al., 2022) and JAQKET  [^7] https://www.nlp.ecei.tohoku.ac.jp/projects/jaqket/), Natural language inference (JNLI (Kurihara et al., 2022)), Summarization (XLSum-ja  [^8] https://huggingface.co/datasets/mkshing/xlsum_ja), Co-reference resolution (WinoGrande-ja (Takahashi et al., 2023)), and Math problems (MGSM (Zhang et al., 2023)).
@@
Due to the unavailability of the MARC-ja dataset (Kurihara et al., 2022), we conducted evaluations on the remaining seven tasks. Given that LLM performance can be significantly influenced by the prompts or templates used, we evaluated all available templates for each task and reported the highest score.
3.2 Compared Models and Instruction Tuning
We compare our pretrained model with six existing Japanese LLMs and three multilingual LLMs. The Japanese LLMs are categorized into two groups: models pretrained from scratch and models that have undergone continual pretraining from multilingual models.